# Machine-learning competition to grade EEG background patterns in newborns with hypoxic-ischaemic encephalopathy

### v1

@@ Abstract
journal: Expert Systems with Applications
@@ 3 Results
Method Weighted MCC Accuracy F1 Precision Recall NEURAL-toolbox 0.761 (0.618–0.909) 0.891 (0.812–0.953) 0.843 (0.710–0.934) 0.873 (0.750–0.963) 0.825 (0.708–0.937) CNN 0.713 (0.566–0.864) 0.859 (0.781–0.937) 0.768 (0.607–0.900) 0.767 (0.621–0.936) 0.813 (0.693–0.915) ConvNeXt 0.677 (0.522–0.835) 0.859 (0.781–0.937) 0.807 (0.677–0.901) 0.862 (0.726–0.965) 0.775 (0.662–0.881) SVM qEEG features 0.437 (0.257–0.627) 0.750 (0.641–0.844) 0.568 (0.328–0.729) 0.821 (0.292–0.879) 0.557 (0.372–0.732)
@@ 4 Discussion
Notably, the best-performing model in our competition utilized only four EEG channels from an 8-channel bipolar montage, reaffirming findings by Raurale et al. (2021) that reduced-channel systems can maintain robust performance [5]. This approach holds promise for improving accessibility in resource-limited settings, where multi-channel EEG systems may not be available. Further validation of reduced-channel configurations is necessary to confirm their clinical utility without sacrificing significant diagnostic accuracy as a reduced channel montage has obvious practical advantages. Additionally, techniques that require fewer EEG channels could be useful for amplitude-integrated EEG (aEEG) systems, the most common form of EEG monitoring available in NICUs, therefore making them available to a broader group of clinicians and research centres.
@@
Additionally, research indicates variability in inter-rater agreement regarding expert classifications of EEG background patterns in newborns with HIE. Wusthoff (2017) [46] reported the highest Cohen-K coefficient of 0.7 (p ¡ 0.001) using a five-grade scale. In contrast, Shavonne (2019)[47] emphasised the necessity for educational and technological strategies to enhance performance in EEG background scoring. These findings reveal a degree of discretion in grading EEG background patterns in newborns with HIE, which affects the ability of a model to generalise to unseen data. This issue can be lessened by training models on extensive and varied datasets labelled by multiple experts. However, adopting a grading standard that minimises human subjectivity might be a more effective approach.

### v2

@@ Abstract
journal:
@@ 3 Results
Method Weighted MCC Accuracy F1 Precision Recall XGBoost NEURAL-toolbox 0.761 (0.618–0.909) 0.891 (0.812–0.953) 0.843 (0.710–0.934) 0.873 (0.750–0.963) 0.825 (0.708–0.937) CNN 0.713 (0.566–0.864) 0.859 (0.781–0.937) 0.768 (0.607–0.900) 0.767 (0.621–0.936) 0.813 (0.693–0.915) ConvNeXt 0.677 (0.522–0.835) 0.859 (0.781–0.937) 0.807 (0.677–0.901) 0.862 (0.726–0.965) 0.775 (0.662–0.881) SVM qEEG features 0.437 (0.257–0.627) 0.750 (0.641–0.844) 0.568 (0.328–0.729) 0.821 (0.292–0.879) 0.557 (0.372–0.732)
@@ 4 Discussion
It must be noted that the scope of this study is not to provide clinical validation of ML models but to test the concept of running ML competitions to crowdsource model development and incentivize data sharing. None of the ML models in this study have been developed to be adopted cot-side by clinicians in real-world scenarios.
The best-performing model in our competition utilized only four EEG channels from an 8-channel bipolar montage, reaffirming findings by Raurale et al. (2021) that reduced-channel systems can maintain robust performance [5]. This approach holds promise for improving accessibility in resource-limited settings, where multi-channel EEG systems may not be available. Further validation of reduced-channel configurations is necessary to confirm their clinical utility without sacrificing significant diagnostic accuracy as a reduced channel montage has obvious practical advantages. Additionally, techniques that require fewer EEG channels could be useful for amplitude-integrated EEG (aEEG) systems, the most common form of EEG monitoring available in NICUs, therefore making them available to a broader group of clinicians and research centres.
@@
Additionally, research indicates variability in inter-rater agreement regarding expert classifications of EEG background patterns in newborns with HIE. Wusthoff (2017) [46] reported the highest Cohen-K coefficient of 0.7 (p << 0.001) using a five-grade scale. In contrast, Shavonne (2019)[47] emphasised the necessity for educational and technological strategies to enhance performance in EEG background scoring. These findings reveal a degree of discretion in grading EEG background patterns in newborns with HIE, which affects the ability of a model to generalise to unseen data. This issue can be lessened by training models on extensive and varied datasets labelled by multiple experts. However, adopting a grading standard that minimises human subjectivity might be a more effective approach.